OpenStudy (anonymous):

how to solve for w (1-wi)/(1-w)=(6-8i+3z-4zi)/(10-5z)

OpenStudy (fibonaccichick666):

Well, I'd first rid myself of those fractions, can you do that?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

then please do

OpenStudy (anonymous):

10-5z-10wi+5wzi=6-8i+3z-4zi-6w+8wi-3zw+4zwi

OpenStudy (anonymous):

-18wi+wzi+6w+3zw=-4+8z-8i-4zi

OpenStudy (fibonaccichick666):

sorry, give me a sec I have my own question up and it's getting answered much quicker than I thought it would

OpenStudy (anonymous):

w=(-4+8z-8i-4zi)/(6+3z-18i+zi)

OpenStudy (anonymous):

how would i simplify this

OpenStudy (fibonaccichick666):

I got slightly different fraction

OpenStudy (fibonaccichick666):

\[\frac{8z-4}{-18i+zi+6+3z}\]

OpenStudy (ikram002p):

\(\frac{1-wi}{1-w}=\frac{6-8i+3z-4zi}{10-5z}\\ {\color{Red} {(1-wi)}}{\color{Magenta} {(10-5z)}}={\color{Magenta} {(6-8i+3z-4zi )}} {\color{Red} {(1-w)}} \\ 10-5z-10wi+5wzi=6-8i+3z-4zi -6w+8wi-3zw+4zwi\\ {\color{Blue} {(10-5z)}}+{\color{Green} {(5wz-10w)}}i={\color{Blue} {(6+3z-6w-3zw)}}+{\color{Green} {(4zw+8w-4z-8)}}i\\ \text{so u would have two equations :-}\\ \text{equation 1 }\\ 10z-5z=6+3z-6w-3zw \text{equation 2 }\\ 5wz-10w=4zw+8w-4z-8\)